how do I evaluate: Log1/5 125?
Answer by lwsshak3(6490)

how do I evaluate: Log1/5 125?
..
Log1/5 125?
Change to exponential form: Base( 1/5) raised to log(x) of number= number(125)
(1/5)^x=125
1/5^x=5^3
5^-x=5^3
-x=3
x=-3
..
ans: The logarithm of 125, base 1/5 = -3